OpenStudy (anonymous):

3/5 of the participants in a triathlon are male, and 1/2 of those are amateurs. What fraction of the racers are amateur males? 6/5 3/10 11/10 4/5

OpenStudy (kropot72):

\[\frac{3}{5}\times \frac{1}{2}\]